# Fact, Fetch, and Reason: A Unified Evaluation of Retrieval-Augmented Generation

### v1

@@ 3.2 Multi-Step Evaluations
Figure 5: This plot shows the accuracy of Gemini-Pro-1.5-0514 on each reasoning type in our test set. We observe a significant increase in performance for all reasoning types when we use multi-step search planning, with the performance on numerical reasoning even exceeding oracle performance.
Multi-Step retrievals significantly improve model performance.
We plot model performance on different combinations of ($k$, $n$, $n\_docs$) in Figure 5. Based on these results, we observe a steady increase in performance as the number of steps and queries are increased with accuracy increasing from $\sim 0.45$ to $\sim 0.52$ for the case of ($k$, $n$, $n\_docs$)=(5,5,2) for vanilla setting where the model is not provided with any specific planning instructions. This is expected, as more steps and queries allow the model to add more relevant documents to its context, leading the model to improve recall, which translates to better accuracy. However, the performance still remains quite low even after five iterations of search retrievals, which is computationally expensive since this requires six non-parallelizable inference calls (five for retrieval + one for final answering) to answer each question. One of the reasons we found behind this slow progress in performance is the lack of diversity in the queries generated by the model; it seemed like the model goes in the wrong direction in search retrievals and never corrects itself. To mitigate this problem, we experiment with two changes to the instructions: (1) We provide a few examples of how an ideal best-case search query sequence should look, and (2) We provide instructions not to repeat queries and force the model to "think step-by-step" (Kojima et al., 2022). We observe a very promising trend with these changes, where the model performance (0.66) through iterations reaches close to the oracle performance (0.73) by the end of five iterations of retrievals. We hope our benchmark will be useful for the community to further reduce the number of search calls and improve model accuracy.
@@ 4 Related Works
Evaluating Retrieval-Augmented Generation (RAG) systems has become increasingly important as these models integrate retrieval mechanisms with generative capabilities to enhance factual accuracy and reasoning(Yu et al., 2024b). Existing benchmarks, such as NaturalQuestions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and ELI5 (Fan et al., 2019), have been used to evaluate RAG models, but they often focus on specific aspects like retrieval accuracy or single-turn question answering without considering the full complexity of real-world applications. For instance, NaturalQuestions primarily tests retrieval precision, while TriviaQA emphasizes factual correctness in trivia-style questions. ELI5, on the other hand, is designed for explainability but does not rigorously assess the multi-hop reasoning necessary for synthesizing information from multiple sources. These benchmarks, while valuable, tend to evaluate RAG systems in a piecemeal fashion, missing the comprehensive assessment needed to truly measure their end-to-end capabilities. We provide additional comparisons against other datasets in Table 1.
FRAMES addresses these limitations by offering a unified and more holistic evaluation framework for RAG systems. Unlike existing datasets, FRAMES tests models across three critical dimensions: factual retrieval, reasoning, and synthesis. It incorporates complex multi-hop queries that require models to retrieve and integrate information from various sources while also handling temporal disambiguation—a challenge not adequately covered by benchmarks like NaturalQuestions or ELI5. Additionally, FRAMES includes tasks that assess the synthesis of information into coherent and contextually accurate responses, ensuring that RAG systems are evaluated on their ability to perform in realistic, multifaceted scenarios. This makes FRAMES a more rigorous and comprehensive benchmark, well suited for guiding the development of next-generation RAG systems.
@@ 5 Conclusion
Future Work.
Moving forward, there are several promising avenues for future research. First, the development of more sophisticated retrieval strategies is essential. This includes exploring dense retrievers trained directly on the multihop retrieval task, such as those based on ColBERT (Khattab & Zaharia, 2020), or SimCSE (Gao et al., 2021) architectures. These approaches could better handle diverse and complex queries by adapting to the context iteratively. Second, improving the reasoning capabilities of LLMs remains a significant challenge. We can explore process supervision methods like those used in PRM-800K (Lightman et al., 2023), or investigate distillation techniques on successful trajectories, similar to approaches in ToolFormer (Schick et al., 2024) and DSPy(Khattab et al., 2023). These methods could enhance numerical, temporal, and post-processing reasoning. Additionally, we can explore modeling approaches such as context reduction of wiki articles to improve planning capabilities and training query generators for more effective information retrieval. Lastly, expanding the FRAMES dataset to include more diverse and domain-specific questions, as well as incorporating more dynamic elements such as real-time information retrieval, could further enhance its utility as a benchmark for next-generation RAG systems. It is important to note that future work should also address the potential limitations of our current approach, including the risk of pretraining data contamination, which may affect the generalizability and reliability of the results, particularly when using Wikipedia articles that could overlap with LLM training data.
Limitations.
@@
Appendix A Synthetic Data Generation Prompt
Synthetic Data Generation System: You are a helpful assistant. User: """TASK: You will be provided with {k_context} Wikipedia article extracts. Based on these extracts, generate {n_questions} challenging factoid questions that meet the following criteria: 1. Standalone & Context-Independent: Questions should not contain any references to "Article 1", "Article 2", etc. They should be understandable without any additional context. 2. Unambiguous Answer: Each question should have a single, clear, and factual answer. 3. Multi-hop Reasoning: Answering each question should require combining information from ALL {k_context} provided Wikipedia articles. 4. Grounded in Context & Conceptual Format: Each question must conceptually follow this format, seamlessly integrating information from each article: **Start with a clear question word (What/How/Where/When).** **Introduce information from each article step-by-step, using connectors to link them logically.** **Example connectors: ’in relation to’, ’compared to’, ’as a result of’, ’which also’, ’in addition to’. ** For each question: * **Provide the single-word answer in parentheses after the question mark.** * **On a new line, clearly explain the reasoning process.** * **For each article, bullet point the specific piece of information used to formulate the question.** Example: **Question:** What type of bird, belonging to the Ardeidae family, went extinct around 1690 and was known for its terrestrial abilities? (Dodo) **Reasoning:** * **Article 1:** Provides information about the Dodo belonging to the Ardeidae family. * **Article 2:** Mentions the extinction of the Dodo around 1690. * **Article 3:** Highlights the Dodo’s adaptation to terrestrial life. {WIKI ARTICLES} """
Figure 6: Prompt used to generate questions synthetically using Gemini-Pro-1.5-0514. k_context and n_questions are placeholders for the number of articles provided and the number of questions to generate per inference.
Appendix B Autorater Prompt
Auto-rating Prompt System: You are a helpful assistant. User: """$=$Task$=$ I need your help in evaluating an answer provided by an LLM against a ground truth answer. Your task is to determine if the ground truth answer is present in the LLM’s response. Please analyze the provided data and make a decision. $=$Instructions$=$ 1. Carefully compare the "Predicted Answer" with the "Ground Truth Answer". 2. Consider the substance of the answers – look for equivalent information or correct answers. Do not focus on exact wording unless the exact wording is crucial to the meaning. 3. Your final decision should be based on whether the meaning and the vital facts of the "Ground Truth Answer" are present in the "Predicted Answer:" $=$Input Data$=$ - Question: <<question>> - Predicted Answer: <<LLM_response>> - Ground Truth Answer: <<ground_truth_answer>>
$=$Output Format$=$ Provide your final evaluation in the following format: "Explanation:" (How you made the decision?) "Decision:" ("TRUE" or "FALSE" ) Please proceed with the evaluation.""" """
Figure 7: Prompt used to auto-rate the responses of LLM in the experiments. The LLM is provided with questions, model responses, and ground truth answers, along with instructions to check if the model response contains the gold answer.

### v2

@@ 3.2 Multi-Step Evaluations
Multi-Step retrieval significantly improves model performance.
We plot model performance on different combinations of ($k$, $n$, $n\_docs$) in Figure 3(b). Based on these results, we observe a steady increase in performance as the number of steps and queries are increased with accuracy increasing from $\sim 0.45$ to $\sim 0.52$ for the case of ($k$, $n$, $n\_docs$)=(5,5,2) for vanilla setting where the model is not provided with any specific planning instructions. This is expected, as more steps and queries allow the model to add more relevant documents to its context, leading the model to improve recall, which translates to better accuracy. However, the performance still remains quite low even after five iterations of search retrievals, which is computationally expensive since this requires six non-parallelizable inference calls (five for retrieval + one for final answering) to answer each question. One of the reasons we found behind this slow progress in performance is the lack of diversity in the queries generated by the model; it seemed like the model goes in the wrong direction in search retrievals and never corrects itself. To mitigate this problem, we experiment with two changes to the instructions: (1) We provide a few examples of how an ideal best-case search query sequence should look, and (2) We provide instructions not to repeat queries and force the model to "think step-by-step" (Kojima et al., 2022). We observe a very promising trend with these changes, where the model performance (0.66) through iterations reaches close to the oracle performance (0.73) by the end of five iterations of retrievals, as shown in Figure 4. We hope our benchmark will be useful for the community to further reduce the number of search calls and improve model accuracy.
@@ 4 Related Works
Evaluating Retrieval-Augmented Generation (RAG) systems is crucial as they combine retrieval and generative capabilities (Yu et al., 2024b). Existing benchmarks like NaturalQuestions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and ELI5 (Fan et al., 2019) focus on specific aspects but lack comprehensive assessment. NaturalQuestions tests retrieval precision, TriviaQA emphasizes factual correctness, and ELI5 targets explainability without rigorous multi-hop reasoning evaluation.
FRAMES addresses these limitations by offering a unified evaluation framework. It tests RAG systems across factual retrieval, reasoning, and synthesis dimensions. Unlike existing datasets, FRAMES incorporates complex multi-hop queries requiring information retrieval and integration from various sources, while handling temporal disambiguation. It also assesses information synthesis into coherent responses, evaluating RAG systems in realistic, multifaceted scenarios. This makes FRAMES a more rigorous and comprehensive benchmark for guiding next-generation RAG system development. Additional dataset comparisons are provided in Table 1.
@@ 5 Conclusion
Limitations
@@
Ethical Considerations
In developing FRAMES, we recognize the importance of addressing potential ethical considerations in retrieval-augmented generation (RAG) research. Researchers using this benchmark should be mindful of possible biases in information sources, the need for safeguards against potential misuse of advanced RAG systems, and the environmental impact of computationally intensive approaches. We encourage exploring efficient, accessible implementations and maintaining transparency in methodologies. By openly discussing these considerations, we aim to foster responsible advancement of RAG capabilities. We welcome ongoing dialogue within the research community to collectively navigate these important ethical aspects as the field progresses.
Acknowledgements
We thank Abhimanyu Goyal, Tsendsuren Munkhdalai, and Dipanjan Das for helpful feedback on writing and presentation of this work.
Appendix A Future Work
Moving forward, there are several promising avenues for future research. First, the development of more sophisticated retrieval strategies is essential. This includes exploring dense retrievers trained directly on the multihop retrieval task, such as those based on ColBERT (Khattab and Zaharia, 2020), or SimCSE (Gao et al., 2021) architectures. These approaches could better handle diverse and complex queries by adapting to the context iteratively. Second, improving the reasoning capabilities of LLMs remains a significant challenge. We can explore process supervision methods like those used in PRM-800K (Lightman et al., 2023), or investigate distillation techniques on successful trajectories, similar to approaches in ToolFormer (Schick et al., 2024) and DSPy (Khattab et al., 2023). These methods could enhance numerical, temporal, and post-processing reasoning. Additionally, we can explore modeling approaches such as context reduction of wiki articles to improve planning capabilities and training query generators for more effective information retrieval. Lastly, expanding the FRAMES dataset to include more diverse and domain-specific questions, as well as incorporating more dynamic elements such as real-time information retrieval, could further enhance its utility as a benchmark for next-generation RAG systems. It is important to note that future work should also address the potential limitations of our current approach, including the risk of pretraining data contamination, which may affect the generalizability and reliability of the results, particularly when using Wikipedia articles that could overlap with LLM training data.